# BEVPlace++: Fast, Robust, and Lightweight LiDAR Global Localization for Unmanned Ground Vehicles

### v2

@@ VI-E Application
We evaluate the loop closing performance of BEVPlace++ and LCDNet by integrating these methods into a state-of-the-art LiDAR SLAM system, i.e. A-LOAM. We conduct our evaluation on sequence “08” of the KITTI dataset, which features challenging reverse loops.
@@ VI-F Understanding BEVPlace++
The design of REM. We conduct experiments to explore the performance of BEVPlace++ with different designs in the feature encoder. We replace the local feature encoder REM with a ResNet34 [26] and test the method on the rotated KITTI dataset to validate the significance of the rotation equivariance design. Additionally, we use different backbone CNNs in REM to study the robustness of BEVPlace++ to various backbones. Tab. X presents the place retrieval and pose estimation results of BEVPlace++ on KITTI. From Tab. X, we can conclude three observations:
@@
  • BEVPlace++ is robust to the choice of CNN backbone in REM. It achieves the best place retrieval recalls and pose estimation success rates when using ResNet34 as the backbone in REM. However, the performance differences when using VGG16 [59], MobileNet [69], and EfficientNet [60] as CNN backbones in REM are not significant.
TABLE X: Place retrieval and pose estimation performance of BEVPlace++ using different feature encoders
@@
Parameter sensitivity. There are two main parameters in the BEVPlace++ network: the number of rotations $N_{R}$ in the REM module and the number of clusters $K$ in NetVLAD. We design two independent experiments on KITTI, with each experiment varying only one parameter, to discover the influence of these parameters. As shown in Tab. XI, the recall rate increases with the number of orientation intervals $N_{R}$ and tends to saturate when $N_{R}\geq 8$. This is reasonable since more accurate rotation equivariance for local features is achieved with larger $N_{R}$. Considering both computational complexity and localization recall, we set $N_{R}=8$. The recall rate increases with the number of NetVLAD clusters $K$, but does not show significant improvement when $K\geq 64$. Therefore, we set $K=64$.
TABLE XI: The recall rates of global localization under different parameter settings of REIN
@@
TABLE XII: Global localization results under different settings of BEV grid
@@
The grid size $g$ is the key parameter to BEV image generation. We conduct experiments on sequences “00” and “08” to evaluate its influence on global localization. As shown in Tab. XII, The success rate decreases and the mean translation and rotation errors tend to increase as the grid size gets large. This is intuitive since the size of the BEV image will get small and the image contents will be highly compressed. We set $g\geq 0.4$ in our experiments by trading between accuracy and computation complexity.
Figure 10: Comparison of detected loops of LCDNet and BEVPlace++. Green lines indicate true positives and red lines indicate false positives. x is the start point of the trajectories.
Figure 11: Visulization of the localization results of sequence 2012-02-25 of NCLT. (a) shows the correct (colored purple) and incorrect localizations (colored red) of LCDNet and BEVPlace++. (b) Zoom-in views of hard scenes for localization. (c) The intermediate localization results of BEVPlace++ including the query BEV, the top-1 candidate of place retrieval, the feature matching correspondence, and the registration results. The query BEV is colored red for better visualization.
@@
Fig. 11 (a) visualizes the results of localizing the point clouds from sequence 2012-02-25 on the global map of NCLT using BEVPlace++ and LCDNet. It demonstrates that BEVPlace++ can successfully perform global localization in more areas than LCDNet. The failed localizations (colored red) of BEVPlace++ primarily occur in challenging scenes such as long corridors and open areas with few measurements. We provide a zoomed-in view of these challenging scenes in Fig. 11 (b) to better illustrate the difficulties in localization. Fig. 11 (c) further shows intermediate results of BEVPlace++ in these hard scenes, including the query BEV, the top-1 candidate for place retrieval, feature matching correspondence, and warp results. In ① and ④ of Fig. 11 (c), BEVPlace++ retrieves false top-1 candidates for the query. In both cases, the query BEV images lack sufficient structural information, hindering BEVPlace++ from extracting global descriptors with enough distinctiveness. In contrast, ② and ③ are examples where BEVPlace++ successfully localizes the queries despite significant view changes.
@@ VII Conclusion
In this paper, we introduce BEVPlace++, a novel global localization method. BEVPlace++ adopts a two-stage method paradigm, sequentially performing place recognition and pose estimation. Utilizing BEV images, BEVPlace++ employs the rotation-equivariant network (REM) to extract robust local features. It generates rotation-invariant global descriptors with NetVLAD pooling. As a complete global localization method, BEVPlace++ can perform multiple localization tasks including place retrieval, loop closing, and global localization. A key insight of BEVPlace++ is that CNNs inherently extract distinctive features from BEV images, as demonstrated through statistical analysis under translation movements. The proposed REM enhances this distinctiveness under rotational transformations. Leveraging these characteristics, BEVPlace++ enables pose estimation for point clouds without explicit pose supervision and adapts well to diverse LiDAR scanners and unknown environments. We conducted experiments across five public datasets, showcasing BEVPlace++’s state-of-the-art performance. Additionally, we integrated BEVPlace++ as a loop closing module in a SLAM system, verifying its effectiveness in handling loop closing tasks. Concise Python APIs of our BEVPlace++ have been open-source to contribute to the robotics community. We hope BEVPlace++ will become a promising new paradigm LiDAR global localization.

### v3

@@ VI-E Application
We evaluate the loop closing performance of BEVPlace++ and LCDNet by integrating these methods into a LiDAR SLAM system, i.e., A-LOAM. We conduct our evaluation on sequence “08” of the KITTI dataset, which contains challenging reverse loops. In the experiments, we generate BEV images from the incoming LiDAR scans. Local and global descriptors are computed by BEVPlace++, which are then associated with their corresponding scan poses from the SLAM system. The BEV database is constructed by combining local descriptors, global descriptors, and poses. Once the robot traverses over 100 meters [43], loop detection is initiated. For every new scan, we compare its global descriptor with descriptors in the existing BEV database. A scan is identified as a loop candidate if the minimum descriptor distance is below a predefined empirical threshold (set to 0.5). Once detecting a valid loop, relative poses between the current scan and the candidate are computed using the local descriptors. Under the plane-motion assumption, the roll, pitch, and vertical (z-axis) offsets between the query and the matched loop scan are assumed to be negligible during loop detection. Thus, we use the estimated 3-DoF relative pose (x, y, yaw) as the initial guess for the ICP alignment between the two point clouds, with the remaining three DoF (z, roll, pitch) set to zero. The ICP refinement produces a full 6-DoF relative transformation. When a loop closure is confirmed, a loop constraint is added to the pose graph optimization framework implemented with Ceres [78]. The pose graph optimization propagates corrections to prior SLAM poses, improving the consistency of the SLAM results.
@@
We conducted additional experiments to evaluate the loop closure performance of BEVPlace++ on a real UGV platform. As shown in Fig. 9, our UGV was equipped with a RoboSense-32 LiDAR sensor, and its ground truth poses were obtained using a high-accuracy RTK-GPS system. A SLAM method, i.e., A-LOAM integrated with BEVPlace++ was deployed on the robot to construct a globally consistent map. During this process, we evaluated both the loop closure effectiveness and runtime performance of BEVPlace++.
As shown in Fig. 9, the odometry trajectory of A-LOAM has a significant offset from the ground truth. However, after integrating our BEVPlace++ into the SLAM framework and performing loop correction, the trajectory aligns much more closely with the ground truth, indicating improved localization accuracy. After loop closing, the absolute trajectory error was reduced from 0.52 m to 0.29 m, demonstrating that the cumulative drift is significantly reduced. For a comparative analysis, we compared BEVPlace++ with other methods on the same data using metrics including average precision (AP), maximum F1 score, maximum recall at 100% precision, mean translation error ($e_{t}$), and mean rotation error ($e_{r}$). As presented in Tab. IX, BEVPlace++ consistently outperformed competing approaches across all metrics, demonstrating its superior generalization capability and effectiveness in real-world scenarios.
The proposed method was implemented on an NVIDIA AGX ORIN embedded system. The descriptor extraction and loop candidate search require about 94.9 ms and 1.2 ms, respectively, while RANSAC pose estimation takes 15.1 ms. These results demonstrate that our method operates in real-time on the embedded system, making it both efficient and practical for real-world applications. Furthermore, by implementing descriptor extraction, candidate retrieval, and pose estimation processes in parallel using ROS, the pipeline achieves a real-time performance frequency of up to 10 Hz. A more detailed visualization video of our experiment is included in the multimedia material [^3] https://youtu.be/4i-ZcqwmMJo?si=TkAp-ymskXTMMjCe.
Figure 9: Comparison of the Odometry trajectory of A-LOAM and the loop trajectory after loop correction using BEVPlace++ on our UGV platform.
TABLE IX: Loop closing performance on the UGV platform.
AP max F1 max recall @ 100% precision $e_{t}$ (m) $e_{r}$ ($\circ$)
LCDNet [17] 0.949 0.867 17.5 0.70 1.59
LCRNet [18] 0.973 0.921 37.8 0.59 1.42
EgoNN [16] 0.922 0.826 57.4 0.54 1.31
BEVPlace++ 0.999 0.982 94.9 0.41 0.98
@@ VI-F Understanding BEVPlace++
The backbone choice of REM. We conduct experiments to explore the performance of BEVPlace++ with different designs in the feature encoder. We replace the local feature encoder REM with a ResNet34 [65] and test the method on the rotated KITTI dataset to validate the significance of the rotation equivariance design. Additionally, we use different backbone CNNs in REM to study the robustness of BEVPlace++ to various backbones. Tab. X presents the place recognition and pose estimation results of BEVPlace++ on KITTI. From Tab. X, we can conclude two observations:
@@
  • BEVPlace++ is robust to the choice of CNN backbone in REM. It achieves the best place recognition recalls and pose estimation success rates when using ResNet34 as the backbone in REM. However, the performance differences when using VGG16 [66], MobileNet [79], and EfficientNet [67] as CNN backbones in REM are not significant.
Different rotation equivariant encoders for BEVPlace++. We also conduct an additional ablation study on the KITTI, NCLT, MCD, and Mulran datasets for comparing different rotation equivariant feature encoders in BEVPlace++, including group convolution [80] used in BEVPlace, a commonly used rotation equivariant library E2CNN [81], and our proposed REM. As shown in Tab. XI, BEVPlace++ achieves the highest recall, success rate of global localization (SR), absolute translation error ($e_{t}$), and absolute rotation error ($e_{r}$) when using REM, highlighting the effectiveness of our designed REM in preserving feature distinctiveness.
TABLE X: Place recognition and pose estimation performance of BEVPlace++ using different feature encoders
@@
TABLE XI: Place recognition and pose estimation performance of BEVPlace++ using different rotation equivariant feature encoders.
KITTI 08 NCLT 2012-02-15 MCD ntu_day_02 Mulran
Recall (%) SR (%) $e_{t}$ (m) $e_{r}$ () Recall (%) SR (%) $e_{t}$ (m) $e_{r}$ () Recall (%) SR (%) $e_{t}$ (m) $e_{r}$ () Recall (%) SR (%) $e_{t}$ (m) $e_{r}$ ()
E2CNN [81] 91.4 94.7 0.67 0.65 67.5 71.4 0.56 1.28 56.9 55.4 0.90 1.53 60.1 61.3 0.75 0.66
Group Conv [80] 92.0 95.7 0.58 0.62 93.5 91.3 0.54 1.21 79.1 75.2 0.83 1.30 80.9 77.8 0.63 0.58
REM 97.3 98.5 0.54 0.57 95.3 95.6 0.32 1.06 83.1 77.9 0.75 1.08 83.1 80.7 0.49 0.40
Parameter sensitivity. There are two main parameters in the BEVPlace++ network: the number of rotations $N_{R}$ in the REM module and the number of clusters $K$ in NetVLAD. We design two independent experiments on KITTI, with each experiment varying only one parameter, to discover the influence of these parameters. As shown in Tab. XII, the recall rate increases with the number of orientation intervals $N_{R}$ and tends to saturate when $N_{R}\geq 8$. This is reasonable since more accurate rotation equivariance for local features is achieved with larger $N_{R}$. Considering both computational complexity and localization recall, we set $N_{R}=8$. The recall rate increases with the number of NetVLAD clusters $K$, but does not show significant improvement when $K\geq 64$. Therefore, we set $K=64$.
TABLE XII: The recall rates of global localization under different parameter settings of REIN
@@
TABLE XIII: Global localization under different settings of BEV grid
@@
The grid size $g$ is the key parameter to BEV image generation. We conduct experiments on sequences “00” and “08” to evaluate the influence of $g$ on global localization. As shown in Tab. XIII, the success rate decreases and the mean translation and rotation errors tend to increase as the grid size gets large. This is intuitive since the size of the BEV image will get small and the image contents will be highly compressed. We set $g\geq 0.4$ in our experiments by trading between accuracy and computation complexity.
@@
Fig. 11 (a) visualizes the results of localizing the point clouds from sequence 2012-02-25 on the global map of NCLT using BEVPlace++ and LCDNet. It demonstrates that BEVPlace++ can successfully perform global localization in more areas than LCDNet. The failed localizations (colored red) of BEVPlace++ primarily occur in challenging scenes such as long corridors and open areas with few measurements. We provide a zoomed-in view of these challenging scenes in Fig. 11 (b) to better illustrate the difficulties in localization. Fig. 11 (c) further shows intermediate results of BEVPlace++ in these hard scenes, including the query BEV, the top-1 candidate for place recognition, feature matching correspondence, and warp results. In ① and ④ of Fig. 11 (c), BEVPlace++ retrieves false top-1 candidates for the query. In both cases, the query BEV images lack sufficient structural information, hindering BEVPlace++ from extracting global descriptors with enough distinctiveness. In contrast, ② and ③ are examples where BEVPlace++ successfully localizes the queries despite significant view changes.
Figure 10: Comparison of detected loops of LCDNet and BEVPlace++ on KITTI and NCLT. The blue line represents the ground truth pose trajectory of the sequence. The red x marks the starting point of the trajectory. Green lines indicate true positive loop matches, connecting the positions of the query and the matched frames, while red lines represent false positives.
Figure 11: Visulization of the localization results of sequence 2012-02-25 of NCLT. (a) shows the correct (colored purple) and incorrect localizations (colored red) of LCDNet and BEVPlace++. (b) Zoom-in views of hard scenes for localization. (c) The intermediate localization results of BEVPlace++, including the query BEV, the top-1 candidate of place recognition, the feature matching correspondence, and the registration results. The query BEV is colored red for better visualization.
@@ VII Conclusion
In this paper, we introduce BEVPlace++, a novel global localization method. BEVPlace++ adopts a two-stage method paradigm, sequentially performing place recognition and pose estimation. Utilizing BEV images, BEVPlace++ employs the rotation-equivariant network (REM) to extract robust local features. It generates rotation-invariant global descriptors with NetVLAD pooling. As a global localization method, BEVPlace++ can perform multiple localization tasks, including place recognition, loop closing, and global localization. A key insight of BEVPlace++ is that CNNs inherently extract distinctive features from BEV images, as demonstrated through statistical analysis under translation movements. The proposed REM enhances this distinctiveness under rotational transformations. Leveraging these characteristics, BEVPlace++ enables pose estimation for point clouds without explicit pose supervision and adapts well to diverse LiDAR scanners and unknown environments. We conducted experiments across seven public datasets, showcasing BEVPlace++’s state-of-the-art performance. Additionally, we integrated BEVPlace++ as a loop closing module in a SLAM system, verifying its effectiveness in handling loop closing tasks in a real UGV. BEVPlace++ has been open-sourced to contribute to the robotics community. We hope BEVPlace++ will become a promising new paradigm for LiDAR global localization.